# SounDiT: Geo-Contextual Soundscape-to-Landscape Generation

### v1

@@ 6.2 Main Results
AudioToken (SD1) 95.210 0.505 0.503 0.419 0.484 0.979
AudioToken (SD2) 165.014 0.502 0.504 0.298 0.455 0.980
GlueGen 64.434 0.498 0.499 0.405 0.481 0.805
SounDiT (Ours) 33.246 0.517 0.592 0.465 0.495 0.788
@@
AudioToken (SD1) 138.785 0.585 0.695 0.364 0.534 0.846
AudioToken (SD2) 118.536 0.637 0.740 0.512 0.724 0.855
GlueGen 195.625 0.584 0.645 0.362 0.582 0.815
SounDiT (Ours) 155.471 0.628 0.744 0.521 0.741 0.757
@@
Sound2Scene 122.460 0.600 0.596 0.287 0.378 0.770
@@
AudioToken (SD1) 163.540 0.532 0.555 0.207 0.311 0.789
AudioToken (SD2) 157.706 0.583 0.587 0.251 0.404 0.803
GlueGen 156.777 0.536 0.547 0.233 0.249 0.839
SounDiT (Ours) 113.111 0.575 0.588 0.336 0.392 0.767
Figure 5 presents landscape images generated by our proposed SounDiT and baseline models across SoundingSVI, Sonic Urban, and two benchmark datasets. The results generated by SounDiT show greater geo-contextual similarity to ground truth images, highlighting potential applications of SounDiT in geography and urban planning.
To evaluate model performance, we adopt general metrics (FID, AIS, IIS) in addition to our proposed geo-contextual PSS framework. As reported in Table 3, SounDiT consistently outperforms baselines on both SoundingSVI and SonicUrban datasets, with notable FID improvements (74$\rightarrow$28 and 49$\rightarrow$33), and achieves superior or competitive results across all metrics. SounDiT consistently ranks highest under PSS, indicating its ability to ensure geographic coherence in generated images. These results validate the effectiveness of PSS in measuring geo-contextual coherence beyond general metrics. The performance drop of SounDiT on the smaller IntoTheWild&Landscape dataset further emphasizes the crucial needs for large-scale training datasets such as SoundingSVI and SonicUrban.
@@
User Study. In addition to these quantitative metrics, we further conducted a user study to assess whether the generated landscape images align with human perceptions of geographic contexts. Specifically, participants were asked to complete two matching tasks: (1) selecting the generated image that best aligned with the given soundscape; and (2) identifying which generated landscape image most closely matched a ground-truth landscape. The average matching accuracy is 86.13%, indicating a strong perceptual alignment between soundscapes and their corresponding generated images. This demonstrates the effectiveness of SounDiT for GeoS2L.
@@ 6.3 Ablations
Table 4: Ablation results. We assess the contribution of key components to the overall performance of SounDiT.
[…]
TA-MoE 28.173 0.652
MHCA 34.108 0.602
(a) SounDiT Block Design
[…]
w/ RAG 0.652 0.578
w/o RAG 0.479 0.514
(b) Geo-contextual RAG
We conducted the ablation study on SoundingSVI. First, we replace the TA-PMoE module with a standard multi-head cross-attention layer, forming SounDiT-MHCA. Second, we disable the RAG module in SounDiT-TA-MoE, yielding SounDiT-TA-MoE w/o RAG. As reported in Table 4(a), SounDiT-TA-MoE achieves a lower FID and a higher $PSS_{element}$, indicating that the expert routing (TA-MoE) is essential for modeling multi-level soundscape and visual features, thereby preserving visual quality and geographic coherence in the generated images. Table 4(b) further shows that incorporating geo-contextual RAG inference improves both element- and scene-level PSS of generated landscape images, highlighting the benefits of integrating geographic contexts during inference.
7 Concluding Remarks
In this work, we present Geo-Contextual Soundscape-to-Landscape (GeoS2L), a practically-informed novel generation problem that synthesizes realistic landscape images from environmental soundscapes. We explicitly integrate geographic domain knowledge into multimodal generative modeling. To support GeoS2L, we construct SoundingSVI and SonicUrban, two large-scale, multi-modal, geo-contextual datasets, capturing diverse sounding and visual environments. We propose SounDiT, a novel diffusion-based framework that captures geographic semantics embedded in sounding environments. We further design a novel geo-contextual evaluation framework, the Place Similarity Score (PSS), which assesses geographic and environmental consistency. Extensive experiments demonstrate that SounDiT consistently outperforms existing baselines and can generate realistic landscape images. The proposed geo-contextual datasets, model, and evaluation framework effectively integrate geographic knowledge into generative modeling, establishing a solid benchmark for the GeoS2L task.
Limitations. While our work introduces a novel problem, GeoS2L, several limitations remain. First, the SoundingSVI dataset relies on geotagged audio from Aporee, which is a crowd-sourced platform that may contain spatial inaccuracies. Future efforts may focus on refining spatial alignment and improving dataset quality through more robust data processing procedures. Second, the current PSS evaluation metrics only consider three practically-informed levels. Future studies could integrate more geographic domain-specific knowledge to support geo-contextually coherent image generation.

### v2

@@ 5 Geo-Contextual Evaluation Suite
AudioToken (SD1) [81] 105.472 0.509 0.572 0.527 0.553 0.797
AudioToken (SD2) [81] 149.473 0.504 0.559 0.249 0.311 0.845
GlueGen [63] 74.250 0.500 0.566 0.400 0.461 0.776
PixArt + MHCA [12] 34.108 0.518 0.578 0.474 0.390 0.743
@@
Sound2Scene [72] 49.681 0.503 0.498 0.332 0.324 0.829
@@
AudioToken (SD1) [81] 95.210 0.505 0.503 0.419 0.484 0.979
AudioToken (SD2) [81] 165.014 0.502 0.504 0.298 0.455 0.980
GlueGen [63] 64.434 0.498 0.499 0.405 0.481 0.805
PixArt + MHCA [12] 41.456 0.517 0.592 0.411 0.396 0.796
@@
6 Experiment
6.1 Implementation Details
Model Architecture and Training. We adopt the pre-trained Variational Autoencoder (VAE) modules from Latent Stable Diffusion [66] as our landscape encoder–decoder pair, which were originally trained on the COCO dataset [49]. For soundscape and scene prompt conditioning, we employ the ImageBind-Huge [23], pre-trained on two million AudioSet clips [21] covering a wide range of environmental, chosen for its accuracy–efficiency balance among several encoders[78]. Our SounDiT is trained using a learning rate of $1\times 10^{-4}$. Experiments are conducted on NVIDIA H100, A100 and A6000 GPUs. During training, we set the soundscape guidance scalar to $\mu=1.0$. For the scene condition, we employ a learnable scaling parameter $\alpha$ initialized to 0 to stabilize early training by starting from an identity mapping. At inference stage, we apply classifier-free guidance with a scale of $4.0$ for both the soundscape and the scene prompt.
Baseline Models. We compare against state of the art audio to image models: CoDi [73], Sound2Scene [72], AudioToken [81], GlueGen [63] and a PixArt [12] variant augmented with multihead cross attention for the soundscape condition (PixArt+MHCA). For CoDi, we utilized its official pre-trained model. For AudioToken, we employed the pre-trained model without additional training, and trained it on our datasets using Stable Diffusion 1 and 2 as its backbone, respectively. The other baselines are training from scratch based on their released code and training settings.
Benchmark Evaluation Metrics. To evaluate the model performance, we adopted FID [33], AIS [33], and IIS [5], effective in previous audio-to-image generation studies [5, 71]. FID compares Gaussian distributions of features extracted from real and generated images, using a pre-trained Inception model. AIS measures audio-image semantic similarity using Wav2CLIP [79]. IIS evaluates image-image semantic similarity using CLIP-based embedding [32].
6.2 Main Results
Figure 3: Visual comparison of landscape images generated by baseline models (CoDi, Sound2Scene, AudioToken(AT), GlueGen, PixArt) and our proposed SounDiT across SoundingSVI and SonicUrban. Ground truth landscape images are provided in the last column. Our proposed SounDiT significant improvements over baseline models in generating geographical context.
[…]
Table 3: Component ablations on SoundingSVI (169K) with two experts in the MoE Soundscape Conditioning module. We assess the individual contributions of Scene Low-Rank Content Mixer (SLRCM) and the S-AdaLN.
[…]
Full Model 19.195 0.538 0.750 0.734
w/o SLRCM + S-AdaLN 25.375 0.511 0.539 0.428
w/o SLRCM 20.335 0.534 0.728 0.704
w/o S-AdaLN 23.435 0.529 0.629 0.572
Practical Applications. To enhance real-world utility and user control, SounDiT supports scene-conditioned generation. Given a fixed soundscape, modifying the scene prompt generates geo-contextual, semantically coherent images that preserve both auditory information and scene context (Fig. 4). This supports practical applications such as soundscape-guided urban design, informing evidence-based design strategies to promote public health, safety, and environmental comfort. First, SounDiT can support soundscape-guided urban design by generating plausible visual place contexts from a target soundscape (quiet/restorative or a recorded clip), helping designers plan actionable elements such as greenery and water. Second, SounDiT facilities noise mitigation strategies by visualizing the built environment, enabling diagnosis of noise sources beyond decibel (dB), informing interventions (green buffers, traffic calming) for planning decisions. Moreover, SounDiT enables inclusive, accessible design by translating soundscape into visual landscapes, broadens participation in urban design/planning practices, especially for hearing-impaired participants.
Figure 4: Scene-conditioned generation. With the same soundscape input, SounDiT generates visually distinct yet acoustically consistent images.
@@ 6.2 Main Results
User Study. In addition to these quantitative metrics, we further conducted a user study to assess whether the generated landscape images align with human perceptions of geographic contexts. Specifically, 17 participants were asked to complete two matching tasks: (1) selecting the generated image that best aligned with the given soundscape; and (2) identifying which generated landscape image most closely matched a ground-truth landscape. The average matching accuracy is 86.13%, indicating a strong perceptual alignment between soundscapes and their corresponding generated images. This demonstrates the effectiveness of SounDiT for GeoS2L.
Table 4: Expert scalability of MoE cross-attention on SoundingSVI (169K). We vary the number of audio experts $M\!\in\!\{2,4,6,8\}$ while keeping other settings fixed.
Experts 2 4 6 8
[…]
$\mathrm{PSS}_{\mathrm{Scene}}\uparrow$ 0.734 0.741 0.742 0.753
@@ 6.3 Ablations
To validate our architectural design and quantify the impact of our proposed modules of SounDiT, we conduct two types of ablation experiments: component ablations and component scalability ablations. First, to assess the contribution of scene prompt and two scene conditioning modules, we perform experiments by first removing both SLRCM and S-AdaLN, and then removing each module individually. As shown in Table 3, removing either module results in a noticeable drop in performance. These results indicate that injecting scene condition both before and after the MoE Soundscape Conditioning module improves geo-contextual alignment and overall quality. Second, we test the scalability of MoE Soundscape Conditioning by varying the number of experts within this module (Table 4). Results show that the model with more experts achieves better performance in FID and scene consistency.
7 Conclusion
In this work, we present Geo-Contextual Soundscape-to-Landscape (GeoS2L), a practically-informed novel generation problem that synthesizes realistic landscape images from environmental soundscapes. To support GeoS2L, we construct SoundingSVI and SonicUrban, two large-scale, multi-modal, geo-contextual datasets, capturing diverse sounding and visual environments. We propose SounDiT, a scalable diffusion-based framework that captures both sounding environment and scene context by MoE Soundscape Conditioning and Scene Low-Rank Content Mixer. We further design a novel geo-contextual evaluation framework, the Place Similarity Score (PSS), which assesses geographic and environmental consistency. Extensive experiments demonstrate that SounDiT consistently outperforms existing baselines and can generate realistic landscape images. The proposed geo-contextual datasets, model, and evaluation framework effectively integrate geographic knowledge into generative modeling, establishing a solid benchmark for the GeoS2L task.
Acknowledgements
We acknowledge ACCESS, supported by the U.S. National Science Foundation, for allocation SOC250002. We also acknowledge resources from the NAIRR Pilot (NAIRR250075). We acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing HPC resources.